OpenStudy (anonymous):

Which expressions are equivalent to the one below? Check all that apply. 10x A. 50x/5x B. 10 • 10x - 1 C. 10 • 10x + 1 D. 50x/5 E. (50/5)x F. x5

OpenStudy (anonymous):

Imagine if the variable x were to be equivalent to 1 or 2. Input that into 10x (giving you 10 or 20) then try that same method on the answer choices and the answer will be the one that matches your 10x :)

OpenStudy (anonymous):

but three are correct

OpenStudy (australopithecus):

It says, Check all that apply. that means there are more than one right answer

OpenStudy (australopithecus):

Which ones do you think are correct?

OpenStudy (anonymous):

a, d, and e?

OpenStudy (australopithecus):

\[\frac{x}{x} = 1\]

OpenStudy (australopithecus):

do you follow?

OpenStudy (anonymous):

no.. i don't get these

OpenStudy (mle):

a number divided by itself would be 1 so x/x would become 1 times what ever else was attached to it and there would be no x left ("a" cough)

OpenStudy (australopithecus):

Well you are mostly right. You seem to get this :) dont be so critical. what I am trying to say is variables divide out. Exponent rules: \[x^{0} = 1\] anything to the power of 0 is equal to 1 \[x^{1} = x\] \[x^{2} = x*x\] \[x^{3} = x*x*x\] etc \[x^{a}x^{b} = x^{a+b}\] \[\frac{x^{a}}{x^{b}} = x^{a}x^{-b} = x^{a + (-b)} = x^{a - b}\] \[x^{-1} = \frac{1}{x^1} = \frac{1}{x}\] \[\frac{1}{x^{-1}} = x^{1} = x\] Examples: \[\frac{34x^2}{x} = 34x\] \[x^5x^3 = x^{5+3} = x^{8}\]

OpenStudy (australopithecus):

more examples: anything raised to the power of 0 is 1: \[(3 + x + 3b)^{0} = 1\] \[\frac{(3+x)}{x+2} = (3+x)(x+2)^{-1}\]

OpenStudy (australopithecus):

does this help?

OpenStudy (anonymous):

It takes me so long to comprehend math

OpenStudy (anonymous):

do the answers have to equal the same thing?

OpenStudy (australopithecus):

yes

OpenStudy (australopithecus):

they all have to equal 10x

OpenStudy (australopithecus):

Look at your answers and then look at the rules I gave you

OpenStudy (australopithecus):

you are on the right track :)

OpenStudy (anonymous):

so i would still go with those answers

OpenStudy (anonymous):

i don't see why I wouldn't?

OpenStudy (preetha):

ThatGirl, please don't forget to rate Australopithecus so he can get his reward!

OpenStudy (mle):

@thatgirlaggie Reconsider choice A 50x/5x can be rewritten as (50/5)*(x/x) as mentioned, (x/x)=1 and since 50/5 = 10 (50/5)*(x/x) would equal 10 (no x)

OpenStudy (australopithecus):

your answers are: \[\frac{50x}{5x}\] \[\frac{50x}{5}\] \[(\frac{50}{5})x\] Multiplication of fraction rule: Note you can divide anything by 1 to make it into a fraction: \[x = \frac{x}{1}\] example: \[62 = \frac{62}{1}\] Multiplication of fractions rule: \[\frac{a}{b}*\frac{b}{c} = \frac{a*b}{b*c}\] example using both \[\frac{52x}{3}*52 = \frac{52x}{3}*\frac{52}{1} = \frac{52x*52}{3*1}\]

OpenStudy (australopithecus):

I am giving you rules so you can do these problems I hope that you take the time to review them, it should help you quite a bit, note a, b an x are all just numbers

OpenStudy (anonymous):

I'm reviewing them and I see what you're saying

OpenStudy (australopithecus):

But yes the answer is D and E \[\frac{50x}{5x} = 10*x^{1-1} = 10*x^{0} = 10 \neq 10x\]

OpenStudy (anonymous):

so because A has two x's, there are only two answers?

OpenStudy (australopithecus):

yes

OpenStudy (anonymous):

i'd say d and e and feel like an idiot now that i get it THANK YOU!!

OpenStudy (australopithecus):

the x cancels out in one of them and you end up with just 10

OpenStudy (mle):

A has x's that cancel out which is why it does Not work :)

OpenStudy (mle):

Don't forget to Rate @Australopithecus for their wonderful help :)

OpenStudy (anonymous):

how do I rate you?

OpenStudy (mle):

top of screen, orange box, "Rate a Qualified Helper"

OpenStudy (australopithecus):

Yeah algebra is pretty simple, I recommend looking up alebra rules and to look at them as you work through these problems. It should help you significantly

OpenStudy (anonymous):

That was wrong

OpenStudy (anonymous):

it was A, C, and E

OpenStudy (australopithecus):

Well they made a mistake lol

OpenStudy (australopithecus):

contact your instructor

OpenStudy (anonymous):

it's a computer. Apex

OpenStudy (mle):

could there have been any typos in your question/options?

OpenStudy (mle):

perhaps the ^2 did not come through a copy paste in 50x^2/5x ?

OpenStudy (australopithecus):

perhaps

OpenStudy (australopithecus):

If that is exactly how it is written and you are being graded by this you should contact their support.

OpenStudy (preetha):

A,C and E can't be right. Check your question. (-:

OpenStudy (anonymous):

what is this section of questions?

OpenStudy (preetha):

Great explanation Australopithecus!

OpenStudy (australopithecus):

Thank you Preetha